The function f(x)=x is uniformly continuous on R but the function f(x)=x^2 is not.Prove
1
Expert's answer
2020-10-29T18:21:54-0400

Definition: A function "f:Dom(f)\to \mathbb{R}" is said to be uniformly continuous if given "\epsilon >0" , there exists "\delta>0" such that whenever "|x-y|<\delta, |f(x)-f(y)|<\epsilon \forall x,y\in Dom(f)" .


We want to show that there exists "\epsilon>0" such that for every "\delta >0 \exist x,y\in\mathbb{R}" such that "|x-y|<\delta" but, "|f(x)-f(y)|\geq\epsilon" . That is, "|x^2-y^2|\geq\epsilon"


Since "f(x)=x^2" ,


"|f(x)-f(y)|=|x^2-y^2|
=|(x-y)(x+y)|\\
=|x-y||x+y|" .


Let "\epsilon=1" , for any "\delta>0" , consider "y=\frac{1}{\delta}" ,


From "|x-y|<\delta,x=\delta+y"


We have,

"|x-y||x+y|=|\delta+y-y||\delta+y+y|\\=\delta\left(\delta+\frac{2}{\delta}\right)=\delta^2+2>1=\epsilon"


"\implies |x^2-y^2|>\epsilon"


Hence, "f(x)=x^2" is not uniformly continuous on "\mathbb{R}" .